# Pinwheel Scheduling with Real Periods

### v1

@@ 1.1 Previous Results
1.1.1 The integer-valued pinwheel scheduling problem.
@@
We will see below that it seems hopeless to extend Theorem 3 directly to the real-valued problem. (Conversely, our proof of Theorem 2 can be seen as a simpler proof of Theorem 3.) The proof of Theorem 3 in the paper [LL97] constructs a valid schedule for a given integer-valued instance $(\underbrace{c_{1},\cdots,c_{1}}_{q_{1}},\underbrace{c_{2},\cdots,c_{2}}_{q_{2}},\underbrace{c_{3},\cdots,c_{3}}_{q_{3}})$. In the schedule, for each $i=1,2,$ and $3$, identify all the tasks with period $c_{i}$ and rename them as task $i$. If the resulting schedule were valid for instance $(\frac{c_{1}}{q_{1}},\frac{c_{2}}{q_{2}},\frac{c_{3}}{q_{3}})$, this would give a solution to the problem with rational periods. However, this attempt fails. We have a counterexample, instance $(\underbrace{13,\cdots,13}_{5},\underbrace{14,\cdots,14}_{5},\underbrace{78,\cdots,78}_{5})$. Apply this to the construction method and identify all the 5 tasks with period 14 in the resulting schedule. It will be found that the identified task is not performed at a period of $\frac{14}{5}$.
@@
If an integer-valued instance $A=(a_{i})_{i\in[k]}$ satisfies $D(A)\leq 1-(a_{1}-1)/(a_{1}(a_{1}+1))$, then $A$ is schedulable.
1.1.2 The real-valued pinwheel scheduling problem.
@@
1.1.3 Other related problems.
@@ 2 A Warm-Up: Case of Two Distinct Real Values
We have found by hand that the three instances each with density 1 in the following Lemma 4 are schedulable and each period is minimum possible.
@@ 3 Proof of Theorem 2
We begin with instances that satisfy the condition of case (I) of Lemma 6, that is $\frac{1}{a_{1}}+\frac{2}{a_{2}}\leq 1$. We here insist on a valid schedule in an explicit form of $S(t)=\ldots$. Otherwise, the schedulability itself can easily be shown by Lemmas 1 and 2, and the schedulability of instance $(a_{1},\frac{a_{2}}{2})$ implied by Theorem 4.
@@
The following lemma ensures the schedulability of six real-valued instances. To find these instances, we formulated a schedule over a finite number of days into a straightforward integer optimization problem (also known as integer program) and solved it using a solver. Each of the periods, except those involving $\varepsilon$, is minimum possible. Substituting zero for $\varepsilon$ makes the instance non-schedulable.
@@ 4 Future Work
Towards Conjecture 2, we would like to build a framework that comprehensively handles real-valued instance sets, analogous to the Pareto surface for integer-valued instance sets [GSW22]. Although the methodology of ours could be extended to the case where the periods take four or five distinct values, solving the general case would be difficult.

### v2

@@ 1.1 Previous Results
The integer-valued pinwheel scheduling problem.
@@
We will see below that it seems hopeless to extend Theorem 3 directly to the real-valued problem. (Conversely, our proof of Theorem 2 can be seen as a simpler proof of Theorem 3.) The proof of Theorem 3 in the paper [LL97] constructs a valid schedule for a given integer-valued instance
$(\underbrace{c_{1},\cdots,c_{1}}_{q_{1}},\underbrace{c_{2},\cdots,c_{2}}_{q_{2}},\underbrace{c_{3},\cdots,c_{3}}_{q_{3}}).$
In the schedule, for each $i=1,2,$ and $3$, identify all the tasks with period $c_{i}$ and rename them as task $i$. If the resulting schedule were valid for instance $(\frac{c_{1}}{q_{1}},\frac{c_{2}}{q_{2}},\frac{c_{3}}{q_{3}})$, this would give a solution to the problem with rational periods. However, this attempt fails. We have a counterexample: instance
$(\underbrace{13,\cdots,13}_{5},\underbrace{14,\cdots,14}_{5},\underbrace{78,\cdots,78}_{5}).$
Apply this to the construction method and identify all the 5 tasks with period 14 in the resulting schedule. It will be found that the identified task is not performed at a period of $\frac{14}{5}$.
@@
If an integer-valued instance $A=(a_{i})_{i\in[k]}$ satisfies $D(A)\leq 1-\frac{a_{1}-1}{a_{1}(a_{1}+1)}$, then $A$ is schedulable.
The real-valued pinwheel scheduling problem.
@@
Other related problems.
@@ 2 A Warm-Up: Case of Two Distinct Real Values
Now we are going to provide some lemmas for the proof of Lemma 3. We have found by hand that the three instances each with density 1 in the following Lemma 4 are schedulable and each period is minimum possible.
@@ 3 Proof of Theorem 2
Lemma 6 splits the whole set of real-valued instances with three periods and density exactly $\frac{5}{6}$ into seven cases (I) through (VII). In particular, for the analysis of cases (II) through (VII), we discovered six real-valued instances of Lemma 8 for which the covering of Lemma 10 succeeds by trial and error while checking schedulability by an integer programming solver.
We discuss some lemmas for the proof of Lemma 6. We begin with instances that satisfy the condition of case (I) of Lemma 6, that is $\frac{1}{a_{1}}+\frac{2}{a_{2}}\leq 1$. We here insist on a valid schedule in an explicit form of $S(t)=\ldots$. Otherwise, the schedulability itself can easily be shown by Lemmas 1 and 2, and the schedulability of instance $(a_{1},\frac{a_{2}}{2})$ implied by Theorem 4.
@@
The following lemma ensures the schedulability of the six real-valued instances. Each of the periods, except those involving $\varepsilon$, is minimum possible. Substituting zero for $\varepsilon$ makes the instance non-schedulable.
@@ 4 Future Work
Towards Conjecture 2, we would like to build a framework that comprehensively handles real-valued instance sets, analogous to the Pareto surface for integer-valued instance sets [GSW22]. Although our current methodology could be extended to the case where the periods take four or five distinct values, solving the general case would be difficult.